# DeepSeekMath: Pushing the Limits of Mathematical Reasoning in Open Language Models

### v1

@@ 1 Introduction
In this study, we introduce DeepSeekMath, a domain-specific language model that significantly outperforms the mathematical capabilities of open-source models and approaches the performance level of GPT-4 on academic benchmarks. To achieve this, we create the DeepSeekMath Corpus, a large-scale high-quality pre-training corpus comprising 120B math tokens. This dataset is extracted from the Common Crawl (CC) using a fastText-based classifier (Joulin et al., 2016). In the initial iteration, the classifier is trained using instances from OpenWebMath (Paster et al., 2023) as positive examples, while incorporating a diverse selection of other web pages to serve as negative examples. Subsequently, we employ the classifier to mine additional positive instances from the CC, which are further refined through human annotation. The classifier is then updated with this enhanced dataset to improve its performance. The evaluation results indicate that the large-scale corpus is of high quality, as our base model DeepSeekMath-Base 7B achieves 64.2% on GSM8K (Cobbe et al., 2021) and 36.2% on the competition-level MATH dataset (Hendrycks et al., 2021), outperforming Minerva 540B (Lewkowycz et al., 2022a). In addition, the DeepSeekMath Corpus is multilingual, so we notice an improvement in Chinese mathematical benchmarks (Wei et al., 2023; Zhong et al., 2023). We believe that our experience in mathematical data processing is a starting point for the research community, and there is significant room for improvement in the future.
@@ 3.2 Training and Evaluating DeepSeekMath-Instruct 7B
  • Open-source models include: general models like (1) DeepSeek-LLM-Chat 67B (DeepSeek-AI, 2024), (2) Qwen 72B (Bai et al., 2023), and (3) ChatGLM3 6B (ChatGLM3 Team, 2023), as well as models with enhancements in mathematics including (4) InternLM2-Math 20B  [^6] https://github.com/InternLM/InternLM-Math which builds on InternLM2 and underwent math training followed by instruction tuning, (5) Math-Shepherd-Mistral 7B which applys PPO training (Schulman et al., 2017) to Mistral 7B (Jiang et al., 2023) with a process-supervised reward model, (6) the WizardMath series (Luo et al., 2023) which improves mathematical reasoning in Mistral 7B and Llama-2 70B (Touvron et al., 2023) using evolve-instruct (i.e., a version of instruction tuning that uses AI-evolved instructions) and PPO training with training problems primarily sourced from GSM8K and MATH, (7) MetaMath 70B (Yu et al., 2023) which is Llama-2 70B fine-tuned on an augmented version of GSM8K and MATH, (8) ToRA 34B Gou et al. (2023) which is CodeLlama 34B fine-tuned to do tool-integrated mathematical reasoning, (9) MAmmoTH 70B (Yue et al., 2023) which is Llama-2 70B instruction-tuned on MathInstruct.
Model Size English Benchmarks Chinese Benchmarks GSM8K MATH MGSM-zh CMATH Chain-of-Thought Reasoning Closed-Source Model Gemini Ultra - 94.4% 53.2% - - GPT-4 - 92.0% 52.9% - 86.0% Inflection-2 - 81.4% 34.8% - - GPT-3.5 - 80.8% 34.1% - 73.8% Gemini Pro - 86.5% 32.6% - - Grok-1 - 62.9% 23.9% - - Baichuan-3 - 88.2% 49.2% - - GLM-4 - 87.6% 47.9% - - Open-Source Model InternLM2-Math 20B 82.6% 37.7% - - Qwen 72B 78.9% 35.2% - - Math-Shepherd-Mistral 7B 84.1% 33.0% - - WizardMath-v1.1 7B 83.2% 33.0% - - DeepSeek-LLM-Chat 67B 84.1% 32.6% 74.0% 80.3% MetaMath 70B 82.3% 26.6% 66.4% 70.9% ChatGLM3 6B 72.3% 25.7% - - WizardMath-v1.0 70B 81.6% 22.7% 64.8% 65.4% DeepSeekMath-Instruct 7B 82.9% 46.8% 73.2% 84.6% DeepSeekMath-RL 7B 88.2% 51.7% 79.6% 88.8% Tool-Integrated Reasoning Closed-Source Model GPT-4 Code Interpreter - 97.0% 69.7% - - Open-Source Model InternLM2-Math 20B 80.7% 54.3% - - DeepSeek-LLM-Chat 67B 86.7% 51.1% 76.4% 85.4% ToRA 34B 80.7% 50.8% 41.2% 53.4% MAmmoTH 70B 72.4% 21.1% - - DeepSeekMath-Instruct 7B 83.7% 57.4% 72.0% 84.3% DeepSeekMath-RL 7B 86.7% 58.8% 78.4% 87.6%
@@ 4.2 Training and Evaluating DeepSeekMath-RL
We conduct RL based on DeepSeekMath-Instruct 7B. The training data of RL are chain-of-thought-format questions related to GSM8K and MATH from the SFT data, which consists of around 144K questions. We exclude other SFT questions to investigate the impact of RL on benchmarks that lack data throughout the RL phase. We construct the training set of reward models following (Wang et al., 2023b). We train our initial reward model based on the DeepSeekMath-Base 7B with a learning rate of 2e-5. For GRPO, we set the learning rate of the policy model as 1e-6. The KL coefficient is 0.04. For each question, we sample $64$ outputs. The max length is set to 1024, and the training batch size is 1024. The policy model only has a single update following each exploration stage. We evaluate DeepSeekMath-RL 7B on benchmarks following DeepSeekMath-Instruct 7B. For DeepSeekMath-RL 7B, GSM8K and MATH with chain-of-thought reasoning can be regarded as in-domain tasks and all the other benchmarks can be regarded as out-of-domain tasks.
@@ 5.2.1 Towards to a Unified Paradigm
There exist three key components: 1) Data Source $\mathcal{D}$, which determines the training data; 2) Reward Function $\pi_{{rf}}$, which is the source of the training reward signal; 3) Algorithm $\mathcal{A}$: which processes the training data and the reward signal to the gradient coefficient $GC$ that determines the magnitude of the penalty or reinforcement for the data. We analyze several representative methods based on such a unified paradigm:

### v2

@@ 1 Introduction
In this study, we introduce DeepSeekMath, a domain-specific language model that significantly outperforms the mathematical capabilities of open-source models and approaches the performance level of GPT-4 on academic benchmarks. To achieve this, we create the DeepSeekMath Corpus, a large-scale high-quality pre-training corpus comprising 120B math tokens. This dataset is extracted from the Common Crawl ($C$) using a fastText-based classifier (Joulin et al., 2016). In the initial iteration, the classifier is trained using instances from OpenWebMath (Paster et al., 2023) as positive examples, while incorporating a diverse selection of other web pages to serve as negative examples. Subsequently, we employ the classifier to mine additional positive instances from the $C$, which are further refined through human annotation. The classifier is then updated with this enhanced dataset to improve its performance. The evaluation results indicate that the large-scale corpus is of high quality, as our base model DeepSeekMath-Base 7B achieves 64.2% on GSM8K (Cobbe et al., 2021) and 36.2% on the competition-level MATH dataset (Hendrycks et al., 2021), outperforming Minerva 540B (Lewkowycz et al., 2022a). In addition, the DeepSeekMath Corpus is multilingual, so we notice an improvement in Chinese mathematical benchmarks (Wei et al., 2023; Zhong et al., 2023). We believe that our experience in mathematical data processing is a starting point for the research community, and there is significant room for improvement in the future.
@@ 3.2 Training and Evaluating DeepSeekMath-Instruct 7B
  • Open-source models include: general models like (1) DeepSeek-LLM-Chat 67B (DeepSeek-AI, 2024), (2) Qwen 72B (Bai et al., 2023), (3) SeaLLM-v2 7B (Nguyen et al., 2023), and (4) ChatGLM3 6B (ChatGLM3 Team, 2023), as well as models with enhancements in mathematics including (5) InternLM2-Math 20B  [^6] https://github.com/InternLM/InternLM-Math which builds on InternLM2 and underwent math training followed by instruction tuning, (6) Math-Shepherd-Mistral 7B which applys PPO training (Schulman et al., 2017) to Mistral 7B (Jiang et al., 2023) with a process-supervised reward model, (7) the WizardMath series (Luo et al., 2023) which improves mathematical reasoning in Mistral 7B and Llama-2 70B (Touvron et al., 2023) using evolve-instruct (i.e., a version of instruction tuning that uses AI-evolved instructions) and PPO training with training problems primarily sourced from GSM8K and MATH, (8) MetaMath 70B (Yu et al., 2023) which is Llama-2 70B fine-tuned on an augmented version of GSM8K and MATH, (9) ToRA 34B Gou et al. (2023) which is CodeLlama 34B fine-tuned to do tool-integrated mathematical reasoning, (10) MAmmoTH 70B (Yue et al., 2023) which is Llama-2 70B instruction-tuned on MathInstruct.
Model Size English Benchmarks Chinese Benchmarks GSM8K MATH MGSM-zh CMATH Chain-of-Thought Reasoning Closed-Source Model Gemini Ultra - 94.4% 53.2% - - GPT-4 - 92.0% 52.9% - 86.0% Inflection-2 - 81.4% 34.8% - - GPT-3.5 - 80.8% 34.1% - 73.8% Gemini Pro - 86.5% 32.6% - - Grok-1 - 62.9% 23.9% - - Baichuan-3 - 88.2% 49.2% - - GLM-4 - 87.6% 47.9% - - Open-Source Model InternLM2-Math 20B 82.6% 37.7% - - Qwen 72B 78.9% 35.2% - - Math-Shepherd-Mistral 7B 84.1% 33.0% - - WizardMath-v1.1 7B 83.2% 33.0% - - DeepSeek-LLM-Chat 67B 84.1% 32.6% 74.0% 80.3% MetaMath 70B 82.3% 26.6% 66.4% 70.9% SeaLLM-v2 7B 78.2% 27.5% 64.8% - ChatGLM3 6B 72.3% 25.7% - - WizardMath-v1.0 70B 81.6% 22.7% 64.8% 65.4% DeepSeekMath-Instruct 7B 82.9% 46.8% 73.2% 84.6% DeepSeekMath-RL 7B 88.2% 51.7% 79.6% 88.8% Tool-Integrated Reasoning Closed-Source Model GPT-4 Code Interpreter - 97.0% 69.7% - - Open-Source Model InternLM2-Math 20B 80.7% 54.3% - - DeepSeek-LLM-Chat 67B 86.7% 51.1% 76.4% 85.4% ToRA 34B 80.7% 50.8% 41.2% 53.4% MAmmoTH 70B 76.9% 41.8% - - DeepSeekMath-Instruct 7B 83.7% 57.4% 72.0% 84.3% DeepSeekMath-RL 7B 86.7% 58.8% 78.4% 87.6%
@@ 4.2 Training and Evaluating DeepSeekMath-RL
We conduct RL based on DeepSeekMath-Instruct 7B. The training data of RL are chain-of-thought-format questions related to GSM8K and MATH from the SFT data, which consists of around 144K questions. We exclude other SFT questions to investigate the impact of RL on benchmarks that lack data throughout the RL phase. We construct the training set of reward models following (Wang et al., 2023b). We train our initial reward model based on the DeepSeekMath-Base 7B with a learning rate of 2e-5. For GRPO, we set the learning rate of the policy model as 1e-6. The KL coefficient is 0.04. For each question, we sample 6464 outputs. The max length is set to 1024, and the training batch size is 1024. The policy model only has a single update following each exploration stage. We evaluate DeepSeekMath-RL 7B on benchmarks following DeepSeekMath-Instruct 7B. For DeepSeekMath-RL 7B, GSM8K and MATH with chain-of-thought reasoning can be regarded as in-domain tasks and all the other benchmarks can be regarded as out-of-domain tasks.
@@ 5.2.1 Towards to a Unified Paradigm
There exist three key components: 1) Data Source $\mathcal{D}$, which determines the training data; 2) Reward Function $\pi_{{rf}}$, which is the source of the training reward signal; 3) Algorithm $\mathcal{A}$: which processes the training data and the reward signal to the gradient coefficient GCGC that determines the magnitude of the penalty or reinforcement for the data. We analyze several representative methods based on such a unified paradigm: